OpenStudy (studiousstudent):

2/3(x+2)-1/4(x-2) Which is an equivalent expression to the one above? (1) 11/12x + 22/12 (2) 11/12x + 10/12 (3) 5/12x + 10/12 (4) 5/12x + 22/12

OpenStudy (phi):

I assume this is \[ \frac{2}{3(x+2)} - \frac{1}{4(x-2)} \]

OpenStudy (phi):

if so, multiply the first fraction by 4(x-2)/4(x-2) (and notice you are doing (a+b)(a-b)= a^2-b^2 ) multiply the 2nd fraction by 3(x+2)/3(x+2)

OpenStudy (phi):

or is it \[ \frac{2}{3}(x+2) - \frac{1}{4}(x-2) \]

OpenStudy (studiousstudent):

Yes phi its the second one

OpenStudy (studiousstudent):

\[2/3(x+2)-1/4(x-2)\]

OpenStudy (phi):

then distribute the 2/3 (that means multiply 2/3 times x and times 2 do the same for -1/4 then you have to "combine" like terms can you do that?

OpenStudy (studiousstudent):

Ill try one minute please

OpenStudy (studiousstudent):

Ill try one minute please

OpenStudy (studiousstudent):

so 2/3x+1.33333-1/4x+0.25

OpenStudy (phi):

yes, but I am guessing they want you to stick with fractions (this question is really about fractions) so (2/3)x + 4/3 - (1/4)x + 2/4 notice that last term is 2/4 (or 1/2) not 0.25 (which means 1/4)

OpenStudy (studiousstudent):

oh I see

OpenStudy (phi):

now combine 2/3 of an x minus 1/4 of an x the common denominator is 12

OpenStudy (studiousstudent):

8/12 and 3/12

OpenStudy (phi):

yes so you now have (8/12) x - (3/12) x +4/3 - 2/4

OpenStudy (phi):

you can now subtract 8/12 x - 3/12 x to get 5/12 x

OpenStudy (studiousstudent):

Isnt it 8/12x+3/12x+4/3+1/2

OpenStudy (studiousstudent):

Or no need to simplify the 2/4

OpenStudy (studiousstudent):

Im confused because dont you have to add to get rid of the -3/12

OpenStudy (studiousstudent):

Wait nevermind I not understand

OpenStudy (phi):

don't lose track of your signs. start at the beginning (2/3)x + 4/3 - (1/4)x + 2/4 you changed the 2/3 to 8/12 and the 1/4 to 3/12 to get (8/12)x + 4/3 - (3/12) x + 2/4 now "collect terms" meaning put the x terms next to each other. notice the -3/12 x is negative

OpenStudy (studiousstudent):

So now we have (5/12) + (4/3) - (1/2)

OpenStudy (studiousstudent):

so like combining them.. pretty much simplifying the expression

OpenStudy (phi):

you mean (5/12)x + (4/3) - (1/2) don't leave out the x now do the same for the numbers, use 12 for the common denominator (because all the answers seem to be using 12)

OpenStudy (studiousstudent):

I have the x on my paper just that I forgot to type it

OpenStudy (studiousstudent):

EDIT: (5/12x) + (16/12) - (6/12)

OpenStudy (phi):

almost done...

OpenStudy (studiousstudent):

(5/12x) + (10/12)?

OpenStudy (phi):

yes. If you are rusty with fractions see http://www.khanacademy.org/math/arithmetic/fractions take a peak at some of the videos.

OpenStudy (studiousstudent):

Yeah i am never rusty with fractions, all these acuity exams I need to improve with them

OpenStudy (studiousstudent):

The acuity exams say thats what I need to improve in

OpenStudy (studiousstudent):

Thank you so much for your help you have earned a fan

OpenStudy (phi):

thanks.

OpenStudy (studiousstudent):

Can you help my out with two more problems ill be posting them now